OpenStudy (anonymous):

Evaluate x^0 + y^0 for x = 3 and y = 2

OpenStudy (anonymous):

2 anything to the 0 is 1 (other than 0^0 = 0)

OpenStudy (anonymous):

anything raised to power zero = 1 (3)^0 + (2)^0 = 1 + 1 = 2